19. To maximize profits in a competitive market set price equal to marginal cost. But in a monopoly, you set marginal eue equal to marginal cost. Based on the given information, answer the following questions. Total Revenue- 100Q-10Q Marginal Revenue = 100-20Q Total Cost 20+10Q Marginal Cost -10 Note: Competitive Market: To maximize profits set Price Marginal Cost Monopoly Market. To maximize profits set Marginal Revenue Marginal Cost a. What price do you charge if you are in a competitive market? b. What quantity do you produce if you are in a competitive market? c. What profits do you make in a competitive market? d. What price do you charge if you are in a monopoly? e. What quantity do you produce if you are in a monopoly? f. What profits do you make if you are in a monopoly? g. Would you rather be in a competitive market or monopoly? Why?
Answer #1

(a)

Total revenue = 100Q - 10Q2

TR = P * Q

P = TR/Q = (100Q - 10Q2)/Q = 100 - 10Q

In order to maximize profit, competitive firm produces that level of output corresponding to which price equals the marginal cost.

P = MC

100 - 10Q = 10

10Q = 90

Q = 9 units

P = 100 - 10Q = 100 - (10*9) = 100 - 90 = 10

Thus,

The price charged in the competitive market is $10 per unit.

(b)

Total revenue = 100Q - 10Q2

TR = P * Q

P = TR/Q = (100Q - 10Q2)/Q = 100 - 10Q

In order to maximize profit, competitive firm produces that level of output corresponding to which price equals the marginal cost.

P = MC

100 - 10Q = 10

10Q = 90

Q = 9 units

Thus,

The quantity produced in the competitive market is 9 units.

(c)

Calculate the profit -

Profit = TR - TC

Profit = (100Q - 10Q2) - (20 + 10Q)

Profit = [(100*9) - (10*92)] - [20 + (10*9)]

Profit = [900 - 810] - [20 + 90]

Profit = 90 - 110

Profit = -20

The profit in a competitive market is -$20.

So, in a sense, there is loss of $20 in the competitive market.

(d)

In order to maximize profit, a monopoly firm produces that level of output corresponding to which MC equals MR.

MR = MC

100 - 20Q = 10

-20Q = -90

Q = 90/20 = 4.5

P = 100 - 10Q = 100 - (100*4.5) = 100 - 45 = 55

Thus,

The price charged in monopoly is $55 per unit.

(e)

In order to maximize profit, a monopoly firm produces that level of output corresponding to which MC equals MR.

MR = MC

100 - 20Q = 10

-20Q = -90

Q = 90/20 = 4.5

Thus,

The quantity produced in the monopoly is 4.5 units.

(f)

Calculate the profit -

Profit = TR - TC

Profit = (100Q - 10Q2) - (20 + 10Q)

Profit = [(100*4.5) - (10*4.52)] - [20 + (10*4.5)]

Profit = [450 - 202.50] - [20 + 45]

Profit = 247.50 - 65

Profit = 182.50

The profit in a monopoly market is $182.50.

(g)

Every firm wants to earn a positive economic profit or at least zero economic profit to remain continue.

It is in the monopoly that firm can earn positive economic profit.

So,

We would rather like to operate in a monopoly.